# Unveiling the Potential of iMarkers: Invisible Fiducial Markers for Advanced Robotics

### v2

@@ Abstract
Fiducial markers are widely used in various robotics tasks, facilitating enhanced navigation, object recognition, and scene understanding. While offering significant advantages for robots and Augmented Reality (AR) applications, they normally disrupt the visual aesthetics of environments because they are visible to humans, making them unsuitable for many everyday use cases. To address this gap, this paper presents iMarkers—innovative, unobtrusive fiducial markers detectable exclusively by robots and AR devices equipped with adequate sensors and detection algorithms. These markers offer high flexibility in production, allowing customization of their visibility range and encoding algorithms to suit various demands. The paper also introduces the hardware designs and software algorithms developed for detecting iMarkers, highlighting their adaptability and robustness in the detection and recognition stages. Various evaluations have demonstrated the effectiveness of iMarkers compared to conventional (printed) and blended fiducial markers and confirmed their applicability in diverse robotics scenarios.
@@ 1 Introduction
Fiducial markers, i.e., artificial landmarks with distinguishable graphical patterns, can bring high potential for various applications where data encoding replaces costly data streaming processes [1]. As markers provide better-defined reference points than naturally available ones, they are simple-yet-effective solutions where reliable feature-matching and pose information extraction are essential. Augmenting information via markers placed in environments and employing proper systems to extract that information can benefit a wide range of cases, including Augmented Reality (AR), Mixed Reality (MR), and robotics tasks [2]. Markers can narrow the range of objects a robot needs to recognize and provide real-time pose information for tracking and localization. For instance, decoding data from markers detected using the robot’s camera is less costly than processing the entire visual scene that the robot can see.
@@
Although such markers can convey various advantages in robotics, seeing multiple objects labeled with printed markers can be aesthetically undesirable and destructive. In other words, many would want to avoid seeing printed tags around them and prefer that the robots operate well without spoiling zones with these artificial landmarks. Moreover, evidence has shown that visible markers can disrupt and interrupt natural gaze behaviors [3]. Consequently, the authors worked on the idea of employing Cholesteric Spherical Reflectors developed by the team to produce a new generation of fiducial markers [4], dubbed iMarker [^1] https://snt-arg.github.io/iMarkers/. This can be done by replacing the ink pigment used in printing common fiducial markers with CSRs of spherical shell type. Due to the peculiar behavior of CSRs, iMarkers can be invisible (indistinguishable) to human eyes but observable by robots in case a dedicated sensor is provided [5]. Importantly, the sensor can be made at a very low cost. Moreover, such markers can be fabricated to be detected only in specific wavelength ranges of light to avoid visual clutter in the environment. The iMarker concept is illustrated in Fig. 1, simulating the effect that the marker is hardly noticeable by the naked eye and in standard camera images while the robot can read and employ them.
@@
  • Introducing iMarkers for robotics and AR/MR; invisible fiducial markers highly flexible in design and undetectable by the naked eye,
  • Proposing multiple sensor solutions and algorithms for detecting iMarkers of various designs and extracting pose information; and
@@ 2.1 Ordinary Fiducial Markers
Considering various fiducial marker libraries available for robotics and AR/MR tasks, they should provide rapid detection, recognition, and information decoding. Ordinary fiducial markers, mainly printed on paper, can be classified into non-square (designed as circles, point sets, or arbitrary visual patterns), square (or matrix-based), and hybrid variations. In non-square markers, identification and data decoding generally occur w.r.t. the center of their geometric shape. InterSense [6] and WhyCode [7] are circular-shaped libraries in which the correspondence point detection is the tricky part. Dot-shaped libraries, such as RUNE-Tag [8], supply fast detection but suffer primarily in noisy conditions. Other arbitrary and shape-free markers are mainly used in 6-Degrees-of-Freedom (6DoF) feature tracking and pose estimation. Despite robust detection, the mentioned method faces challenges like inter-marker encoded dot confusion and steep viewing angle recognition.
Square fiducial markers, on the other hand, employ binary codes encoded in matrices and deliver four principal corner points for more precise pose estimation. They require less computation cost for detection and recognition, leading to potentially faster detection algorithms. ARToolkit [9] is an open-source matrix-based marker library that uses template-matching for square marker pattern determination. However, it may face high re-projection errors with rotation and camera angle changes. AprilTag [10] is another open-source library with robust pose estimation and better false positive confusion rate thanks to its high-speed line/border detection and graph-based image segmentation. Similarly, Garrido-Jurad et al. introduced ArUco Marker [11] that performs well in occlusion and Gaussian noise conditions by efficiently storing and retrieving binary vectors forming marker patterns.
@@ 2.3 Gap Identification
Detection Algorithm Performance: Since the sensors are designed to generate binary images, fiducial marker patterns are detected within milliseconds, ensuring real-time performance for various robotics and AR/MR use cases.
@@ 4.1 Dual-vision Solution
    $f_{p}\leftarrow$ post-process $f_{p}$erosion $+$ Gaussian blur}
@@ 4.2 Dynamic Single-vision Setup
    $f_{p}\leftarrow$ post-process $f_{p}$erosion $+$ Gaussian blur}
@@ 4.3 Static Single-vision Setup
0: frame-set $F$ from the camera ($RGB$)
@@
$r\leftarrow(low,high)$demanded color range in $HSV$}
@@
    $f_{p}\leftarrow$ convert $f$ to color space $HSV$
@@
    $f_{p}\leftarrow$ post-process $f_{p}$erosion $+$ Gaussian blur}
@@
0: frame-set $F$ from the camera ($grayscale$)
@@
$r\leftarrow(low,high)$demanded gray-level range in $grayscale$}
@@
    $f_{p}\leftarrow$ post-process $f_{p}$erosion $+$ Gaussian blur}
@@ 5.1 Operational Implementation
Software. To support the detection algorithms outlined in Section 4, an open-source iMarker detection software is developed, publicly available at https://snt-arg.github.io/iMarkers/. The framework is implemented in Python and features a Robot Operating System (ROS) interface, enabling straightforward integration into robotic applications.
@@ 5.2 Qualitative Evaluation
Given the complexity of measuring aesthetic perception, this section aims to visually assess iMarkers and showcase their detection performance. To achieve this, various iMarkers were strategically placed in rigid real-world locations and captured using different hardware sensors, highlighting their seamless integration into diverse settings. Fig. 7 depicts several qualitative results, showing the versatility of iMarkers across different spectral ranges and detection setups.
@@
Figure 7: iMarker variants detected using different sensor setups and corresponding algorithms, including the visible-range iMarker detection procedures using the static single-vision (row a), dual-vision (row b), and dynamic single-vision (row c) setups; and UV-range iMarker detection procedure using the single-vision setup (row d).
Accordingly, row a of Fig. 7 presents a visible-range green iMarker on a same-color background. In this scenario, Fig. 7-a-i shows its appearance to the naked eye, confirming its transparency and unobtrusiveness. Fig. 7-a-ii depicts its captured shot using the static single-vision setup, and Fig. 7-a-iii demonstrates the detection and pose estimation outcome using Algorithm 3. Row b focuses on the dual-vision setup output aiming at a visible-range iMarker on a patterned surface under low-light conditions. In this case, Figures 7-b-i and 7-b-ii capture images via left- and right-handed polarizers attached to the sensor cameras, respectively. The outcome is shown in Fig. 7-b-iii, where the iMarker is recognized using Algorithm 1. The performance of the dynamic single-vision setup is shown in the row c,” where $f_{t-1}^{\text{~{}off}}$ stores the scene in Fig. 7-c-i and $f_{t}^{\text{~{}on}}$ captures Fig. 7-c-ii. The “temporal” subtraction output based on Algorithm 2 is illustrated in Fig. 7-c-iii. Finally, the UV-range iMarker placed on a laptop guarantees exceptional invisibility to the naked eye (Fig. 7-d-i). However, the inner patterns become visible using a near-UV illumination when captured by a UV-range camera (Fig. 7-d-ii). Applying the methodology introduced in Algorithm 4, as shown in Fig. 7-d-iii, will facilitate the recognition process.
5.3 Detection Range Evaluation
@@
A comparative experiment was conducted using printed markers and iMarkers, formatted as ArUco library [11], to evaluate the effectiveness of iMarkers w.r.t. their classic counterparts. The markers were affixed to a tripod with an adjustable mounting plate, enabling precise control over the viewing angle relative to the camera. The objective was to measure the maximum distance across different viewing angles where each marker variant remained detectable using the ArUco detection algorithm. Accordingly, an Intel® RealSense™ D435 camera was used across all experiments: unmodified for printed markers and equipped with static or dynamic polarizers for iMarker detection, ensuring consistent sensor characteristics for fair comparison. The examined viewing angle ranged from $0^{\circ}$ to $75^{\circ}$, within which the estimated poses of the markers remained reliable. Experiments were conducted under normal lighting conditions, and $7\times 7~{}\mathrm{cm^{2}}$ visible-range iMarker variants were used to be compared against $6\times 6~{}\mathrm{cm^{2}}$ and $7\times 7~{}\mathrm{cm^{2}}$ printed ArUco markers as references. Experimental results are depicted in Fig. 8.
@@
According to the evaluation results, increasing the viewing angle reduces the maximum marker detection range in all scenarios, mainly due to perspective distortion. Additionally, iMarker detection using static and dynamic single-vision setups shows negligible differences in maximum detection ranges, differing between $3~{}\mathrm{cm}$ to $7~{}\mathrm{cm}$, with an average of $4.8~{}\mathrm{cm}$. However, regardless of the detection setup, printed markers achieved consistently higher detection ranges than iMarkers of identical dimensions ($7\times 7~{}\mathrm{cm^{2}}$). The detection range of $7\times 7~{}\mathrm{cm^{2}}$ iMarkers closely matches that of $6\times 6~{}\mathrm{cm^{2}}$ printed ArUco markers, with differences ranging from $4~{}\mathrm{cm}$ to $6~{}\mathrm{cm}$ and averaging $4.2~{}\mathrm{cm}$ across various angles of view. It should be noted that printed ArUco markers, thanks to their high-contrast black-and-white patterns, offer substantial visual features that facilitate detection. However, iMarkers demonstrate comparable detection performance in real-world experiments, apart from their distinct advantage over classic markers: their ability to seamlessly blend into the environment without disrupting visual aesthetics. This makes iMarkers highly suitable for shorter-range applications where visual discretion is essential, such as indoor navigation.
@@ 5.4 Recognition Speed
Considering the necessity of fast recognition speeds in demonstrating the practicality of fiducial markers for real-time applications, this subsection provides a detailed profiling analysis of iMarker detection and recognition. Table 2 summarizes cumulative processing times for iMarker data decoding. Evaluations were conducted using various sensor configurations and implemented software introduced in subsection 5.1. The algorithms developed for all sensor variants share five core modules: acquisition,” which captures video frames; preprocessing,” which applies image enhancement techniques; processing,” which refines frames to emphasize marker patterns; post-processing,” which enhances the final binary images; and recognition to detect and interpret marker patterns. To obtain the measurements, $7\times 7~{}\mathrm{cm^{2}}$ iMarkers were placed approximately $50~{}\mathrm{cm}$ from the lens, directly facing the sensor under standard lighting conditions with evenly diffused artificial illumination. For the UV-range iMarker, a DARKBEAM A300 $365~{}nm$ flashlight was additionally used to enhance the iMarker visibility.

### v3

@@ Abstract
Fiducial markers are widely used in various robotics tasks, facilitating enhanced navigation, object recognition, and scene understanding. While offering significant advantages for robots and Augmented Reality (AR) applications, they normally disrupt the visual aesthetics of environments because they are visible to humans, making them unsuitable for many everyday use cases. To address this gap, this paper presents iMarkers—innovative, unobtrusive fiducial markers detectable exclusively by robots and AR devices equipped with adequate sensors and detection algorithms. These markers offer high flexibility in production, allowing customization of their visibility range and encoding algorithms to suit various demands. The paper also introduces the hardware designs and open-sourced software algorithms developed for detecting iMarkers, highlighting their adaptability and robustness in the detection and recognition stages. Various evaluations have demonstrated the effectiveness of iMarkers compared to conventional (printed) and blended fiducial markers and confirmed their applicability in diverse robotics scenarios.
@@ 1 Introduction
Fiducial markers, i.e., artificial landmarks with distinguishable graphical patterns, are widely used in robotics and computer vision applications to offer reliable object identification, localization, and pose estimation [1]. As markers provide better-defined reference points than naturally available ones, they are simple-yet-effective solutions where consistent feature-matching and precise pose retrieval are essential. Additionally, augmenting information via markers placed in environments and employing proper systems to extract that information can benefit a wide range of cases, including Augmented Reality (AR) and robotics tasks [2]. Markers can narrow the range of objects a robot needs to recognize and provide real-time pose information for tracking and localization. For instance, decoding data from markers detected using the robot’s camera is less costly than processing the entire visual scene that the robot can see.
@@
Although such markers can convey various advantages in robotics, seeing multiple objects labeled with printed markers can be aesthetically undesirable and destructive. In other words, many would want to avoid seeing printed tags around them and prefer that the robots operate well without spoiling zones with these artificial landmarks. Moreover, evidence has shown that visible markers can disrupt and interrupt natural gaze behaviors [3]. Consequently, the authors worked on the idea of employing Cholesteric Spherical Reflectors developed by the team to produce a new generation of fiducial markers [4], dubbed iMarker [^1] https://snt-arg.github.io/iMarkers/. This can be done by replacing the ink pigment used in printing common fiducial markers with CSRs of spherical shell type. Due to the peculiar behavior of CSRs, iMarkers can be invisible (indistinguishable) to human eyes but observable by robots in case a dedicated sensor is provided [5]. Importantly, the sensor can be made at a very low cost. Moreover, such markers can be fabricated to be detected only within specific wavelength ranges of light, thereby avoiding visual clutter in the environment. The iMarker concept is illustrated in Fig. 1, confirming, through qualitative comparison, that they are hardly noticeable by the naked eye and in standard camera images compared to traditional printed fiducial markers.
@@
  • Introducing iMarkers for robotics and AR applications; invisible fiducial markers highly flexible in design and undetectable by the naked eye,
  • Proposing multiple sensor solutions and software algorithms for detecting iMarkers of various designs and extracting their 6DoF pose information; and
@@ 2.1 Ordinary Fiducial Markers
Fiducial marker libraries used in robotics and AR applications are expected to enable rapid detection, reliable recognition, and efficient information decoding. Ordinary fiducial markers, mainly printed on paper, can be classified into non-square (designed as circles, point sets, or arbitrary visual patterns), square (or matrix-based), and hybrid variations. In non-square markers, identification and data decoding generally occur w.r.t. the center of their geometric shape. InterSense [6] and WhyCode [7] are circular-shaped libraries in which the correspondence point detection is the tricky part. Dot-shaped libraries, such as RUNE-Tag [8], supply fast detection but suffer primarily in noisy conditions. Other arbitrary and shape-free markers are mainly used in 6-Degrees-of-Freedom (6DoF) feature tracking and pose estimation. Despite robust detection, the mentioned method faces challenges like inter-marker encoded dot confusion and steep viewing angle recognition.
Square fiducial markers, on the other hand, employ binary codes encoded in matrices and deliver four principal corner points for more precise pose estimation. They require less computation cost for detection and recognition, leading to potentially faster detection algorithms. ARToolkit [9] is an open-source matrix-based marker library that uses template-matching for square marker pattern determination. However, it may face high re-projection errors with rotation and camera angle changes. AprilTag [10] is another marker library that incorporates gradient-based edge detection and reliable line identification to ensure accurate marker recognition, even under challenging viewing angles. Its enhanced digital coding system has made it more robust to rotation, occlusion, and warping compared to traditional systems, such as ARToolkit. Similarly, Garrido-Jurad et al. introduced ArUco Marker [11] that generates configurable marker dictionaries with maximized inter-marker distances. Their system contains automatic marker detection with error correction based on dictionary search and a novel occlusion handling method using color-based occlusion masks for robust pose estimation.
@@ 2.3 Gap Identification
Detection Algorithm Performance: Since the sensors are designed to generate binary images, fiducial marker patterns are detected within milliseconds, ensuring real-time performance for various robotics and AR use cases.
@@ 4.1 Dual-vision Solution
  $f_{p}\leftarrow$ post-process $f_{p}$ {erosion ++ Gaussian blur}
@@ 4.2 Dynamic Single-vision Setup
  $f_{p}\leftarrow$ post-process $f_{p}$ {erosion ++ Gaussian blur}
@@ 4.3 Static Single-vision Setup
0: frame-set $F$ from the camera (RGBRGB)
@@
$r\leftarrow(low,high)$ {demanded color range in HSVHSV}
@@
  $f_{p}\leftarrow$ convert $f$ to color space HSVHSV
@@
  $f_{p}\leftarrow$ post-process $f_{p}$ {erosion ++ Gaussian blur}
@@
0: frame-set $F$ from the camera (grayscalegrayscale)
@@
$r\leftarrow(low,high)$ {demanded gray-level range in grayscalegrayscale}
@@
  $f_{p}\leftarrow$ post-process $f_{p}$ {erosion ++ Gaussian blur}
@@ 5.1 Operational Implementation
Software. To support the detection algorithms outlined in Section 4, an open-sourced  [^2] https://snt-arg.github.io/iMarkers/ iMarker detection software framework has been developed. The framework is implemented in Python and features a $\mathrm{ROS2}$ interface, enabling straightforward integration into robotic applications.
@@ 5.2 Qualitative Evaluation
Given the complexity of measuring aesthetic perception, this section aims to visually assess iMarkers and showcase their detection performance. To achieve this, various iMarkers were strategically placed in rigid real-world locations and captured using different hardware sensors, highlighting their seamless integration into diverse settings. It should be noted that while aesthetic perception is inherently subjective, the visual comparisons provided in Fig. 1 and Fig. 3 offer practical, real-world evidence supporting the unobtrusiveness of iMarkers in diverse environments.
((a)) Static single-vision setup detecting a visible-range iMarker: raw frame (left), polarizer applied (middle), and detection result using color masking (right).
((b)) Dynamic single-vision setup detecting a visible-range iMarker: frames $f_{t-1}^{\text{~off}}$ (left) and $f_{t}^{\text{~on}}$ (middle) captured at different times with alternating polarizer states, and the detection result obtained via temporal subtraction (right).
((c)) Dynamic single-vision setup detecting a visible-range iMarker in an outdoor environment. The order of the frames (off-polarized, on-polarized, and subtraction) follows the same structure as (b).
((d)) Dual-vision setup capturing a visible-range iMarker under low-light conditions: frames $f_{l}$ (left) and $f_{r}$ (middle) acquired using orthogonal polarizers, and the detection result obtained via spatial frame subtraction $f_{l}-f_{r}$ (right).
((e)) Single-vision UV iMarker detection: scene as captured by a standard mobile phone camera (left), corresponding view captured with a UV-sensitive camera under near-UV illumination (middle), and the resulting marker pattern revealing output via thresholding (right).
@@
Figure 8: iMarker variants detected using different sensor setups and corresponding algorithms, including the visible- and UV-range iMarkers.
Fig. 8 depicts several qualitative results, showing the versatility of iMarkers across different spectral ranges and detection setups. Accordingly, Fig. 8(a) presents a visible-range green iMarker discreetly embedded on a similar color background, demonstrating its visual camouflage and unobtrusiveness. In this scenario, the static single-vision setup was employed to capture the scene, where the setup’s polarizer effectively blocks the CSR regions forming the inner iMarker pattern, enhancing the contrast and enabling enhanced separation for detection. The detection and pose estimation outcome was obtained using Algorithm 3. Fig. 8(b) and Fig. 8(c) depict the output of the dynamic single-vision setups applied to the same visible-range iMarker in indoor and outdoor conditions. Detection is achieved through temporal subtraction between frames $f_{t-1}^{\text{~off}}$ and $f_{t}^{\text{~on}}$ following the procedure described in Algorithm 2. In Fig. 8(d), the output of the dual-vision setup targeting a visible-range iMarker placed on a patterned surface under low-light conditions is shown. In this configuration, both cameras simultaneously capture the same scene, with the left camera acquiring frame $f_{l}$ through a left-handed polarizer and the right camera capturing frame $f_{r}$ through a right-handed polarizer. Marker detection was performed using Algorithm 1 , which applies spatial subtraction between the polarized image pairs. Finally, the UV-range iMarker placed on a transparent surface depicts exceptional invisibility to the naked eye, as illustrated in Fig. 8(e). To reveal the inner patterns of the iMarker, a near-UV illumination was applied, and the same scene was captured using the UV-range camera, as shown in the middle sub-figure. The methodology introduced in Algorithm 4 facilitates the marker detection process by applying a tailored thresholding strategy to the UV image, enabling reliable iMarker’s invisible patterns extraction.
5.3 Detection Range and Pose Accuracy Evaluation
@@
To evaluate the effectiveness of iMarkers against their classic paper-printed counterparts, a series of comparative experiments were conducted using both marker variants in the ArUco format. The experiments focused on two primary objectives: (1) determining the maximum detection range across varying viewing angles, and (2) assessing the accuracy of the estimated 6DoF pose for each marker type; both were done using the ArUco detection algorithm. The employed sensor across all scenarios was the single-vision setup with Intel® RealSense™ D435, unmodified for printed markers and equipped with static/dynamic polarizers for iMarker detection. We intentionally limited the analysis to this setup to ensure fair comparison for detecting printed markers and iMarkers, while other sensors in Fig. 6 differ significantly in resolution, shutter type, and optical characteristics. All experiments were conducted using the camera’s RGB sensor at a resolution of $1024\times 768$ to maintain uniform input quality across evaluations. Regarding marker dimensions, $7\times 7~\mathrm{cm^{2}}$ visible-range iMarker variants were used to be compared against $6\times 6~\mathrm{cm^{2}}$ and $7\times 7~\mathrm{cm^{2}}$ printed ArUco marker references. Evaluations were performed under normal lighting conditions and in a controlled robotics laboratory equipped with an OptiTrack motion capture system, providing sub-millimeter ground truth accuracy for 6DoF pose tracking. Both the printed fiducial markers and the fabricated iMarkers were affixed to a tripod with an adjustable mounting plate, allowing precise control over their orientation relative to the camera. To ensure accurate ground truth acquisition during the evaluations, omnireflective motion capture landmarks were placed on both the fiducial markers, with their poses continuously tracked using the OptiTrack system. The examined viewing angle ranged from $0^{\circ}$ to $75^{\circ}$, within which the estimated poses of the markers remained reliable w.r.t. the ground truth.
Maximum detection range. According to the evaluation results in Fig. 9, increasing the viewing angle reduces the maximum marker detection range in all scenarios, mainly due to perspective distortion. Additionally, iMarker detection using static and dynamic single-vision setups shows negligible differences in maximum detection ranges, differing between $3~\mathrm{cm}$ to $7~\mathrm{cm}$, with an average of $4.82~\mathrm{cm}$. However, regardless of the detection setup, printed markers achieved consistently higher detection ranges than iMarkers of identical dimensions ($7\times 7~\mathrm{cm^{2}}$). The detection range of $7\times 7~\mathrm{cm^{2}}$ iMarkers closely matches that of $6\times 6~\mathrm{cm^{2}}$ printed ArUco markers, with differences ranging from $4.12~\mathrm{cm}$ to $6.39~\mathrm{cm}$ and averaging $4.24~\mathrm{cm}$ across various angles of view. It should be noted that printed ArUco markers, thanks to their high-contrast black-and-white patterns, offer substantial visual features that facilitate detection. However, iMarkers demonstrate comparable detection performance in real-world experiments, apart from their distinct advantage over classic markers: their ability to seamlessly blend into the environment without disrupting visual aesthetics. This makes iMarkers highly suitable for shorter-range applications where visual discretion is essential, such as indoor navigation.
@@
Pose accuracy. The pose estimation accuracy was assessed by measuring the deviation of the detected marker poses from ground truth data under the mentioned viewing angles. As shown in Table LABEL:tbl_pose_error, the mean pose estimation error norm across various viewing angles remains consistent among the printed ArUco markers and their iMarker counterparts. In general, the detection error increases as the viewing angle grows from $0^{\circ}$ to $75^{\circ}$ due to increasing perspective distortion of the marker appearance. The relative difference in pose estimation error between printed markers and iMarkers remains within a narrow range of $1.3\%$ to $3.9\%$. This consistency and negligible error differences were expected outcomes of our system design, as the proposed hardware-software configurations are explicitly designed to produce pre-processed binary images revealing the marker patterns prior to the detection and pose estimation stage in the standard ArUco detector library. Thus, the marker detector library can recognize both types of markers by applying built-in preprocessing to either standard RGB inputs of printed markers, or the binary outputs generated by the iMarker detection system.
@@ 5.4 Recognition Speed
Considering the necessity of fast recognition speeds in demonstrating the practicality of fiducial markers for real-time applications, this subsection provides a detailed profiling analysis of iMarker detection and recognition. Table 2 summarizes cumulative processing times for iMarker data decoding. Evaluations were conducted using various sensor configurations and implemented software introduced in subsection 5.1. The algorithms developed for all sensor variants share five core modules: acquisition,” which captures video frames; preprocessing,” which applies image enhancement techniques; processing,” which refines frames to emphasize marker patterns; post-processing,” which enhances the final binary images; and recognition to detect and interpret marker patterns. To obtain the measurements, $7\times 7~\mathrm{cm^{2}}$ iMarkers were placed approximately $50~\mathrm{cm}$ from the lens, directly facing the sensor under standard lighting conditions with evenly diffused artificial illumination. For the UV-range iMarker, a DARKBEAM A300 365nm365~nm flashlight was additionally used to enhance the iMarker visibility.